how do you do equations like... for example -3 tenths h = 15 ? or -1.1k = -1.21
Answer by stanbon(75887)

-3 tenths h = 15
(-3/10)h = 15
-3h = 150
h = -50
--------------------
-1.1k = -1.21
1.1k = 1.21
k = 1.1
==============
Cheers,
Stan H.